OpenStudy (anonymous):

PLEASE HELP ME I WILL MEDAL AND FAN!!!!!!!!!!!! Find the product -320times7/16 A. 280 B. 140 C. –140 D. –280

OpenStudy (lizz123):

\[-321\times \frac{ 7 }{ 16 }\]

OpenStudy (lizz123):

Right under -321 put a one

OpenStudy (lizz123):

 multiply straight across

OpenStudy (anonymous):

okay lol look I don't mean to get into this conversation but I really need help

OpenStudy (lizz123):

wht is -321 * 7 = ?

OpenStudy (lizz123):

@lol23

OpenStudy (lizz123):

It equals-2247. Now what does 1 * 16 = ?

OpenStudy (anonymous):

its 16

OpenStudy (lizz123):

\[\frac{ -2247 }{ 16 }\]

OpenStudy (lizz123):

Simplify that

OpenStudy (lizz123):

The answer is -140 7/16

OpenStudy (lizz123):

So I would go for C. -140

OpenStudy (anonymous):

k tnx got more ill tag u

OpenStudy (lizz123):

k

OpenStudy (anonymous):

now can I get help please